OpenStudy (anonymous):

Write the quadratic function in vertex form. f(x) = -4x2 - 2x -3

OpenStudy (anonymous):

first coordinate of vertex is \(-\frac{b}{2a}\) in this case you have \(-\frac{-2}{2\times -4}=-\frac{1}{4}\) so it will look like \[f(x)=-4(x+\frac{1}{4})^2+k\] to find \(k\) replace \(x\) by \(\frac{1}{4}\) in the the function and see what you get